OpenStudy (anonymous):

Math Analysis: Find the indicated roots. Write the answer in a + bi from Cube roots of 27i

OpenStudy (blockcolder):

Convert to polar, use formula, convert back to rectangular.

OpenStudy (unklerhaukus):

\[\sqrt[3]{27i}=\sqrt[3]{3^3i}=3\sqrt[3]i\]

OpenStudy (anonymous):

3a=90degrees correct? o-o

OpenStudy (anonymous):

is this right..? 3cis(30) 2.60 + 1.5i 3 cis(150) -2.60 + 1.5i 3cis(270) 0 + -3i

OpenStudy (anonymous):

was i supposed to put it in fractions cause i put it in decimals.. don't know if i would be marked wrong for that one :\

OpenStudy (anonymous):

nvm, i got it :p